OpenStudy (anonymous):

let f (x) x2 1 and g x 3x – 1. How would you Find f(g(x))

OpenStudy (campbell_st):

f(g(x)) = f(3x -1) f(3x -1) = (3x -1)^2 + 1 just expand and simplify

OpenStudy (anonymous):

Expand?

OpenStudy (anonymous):

Yes now expand..

OpenStudy (campbell_st):

multiply to remove the bracket... perhaps distribute is you term

OpenStudy (anonymous):

Ummm..... Ill try... I am not sure about it though

OpenStudy (campbell_st):

can you expand a perfect square..?

OpenStudy (anonymous):

\[(a-b)^2 = a^2 + b^2 - 2ab\]

OpenStudy (anonymous):

I have never heard of that before.. to expand it ...

OpenStudy (campbell_st):

ummm thats interesting given the question you are attempting to answer...

OpenStudy (campbell_st):

are the 2 functions \[f(x) = x^2 + 1\] and \[g(x) = 3x - 1\]

OpenStudy (anonymous):

.... Maybe my teacher used a diffrent word? I am not sure.... lol This is why I am lost

OpenStudy (campbell_st):

have you done any work on quadratics...?

OpenStudy (anonymous):

What am I doing to that? and no..... I am taking a quiz and this pops up...and i was dumbfounded

OpenStudy (anonymous):

thats not the right word. clueless is

OpenStudy (campbell_st):

lol... well just use the answer f(g(x)) = (3x -1)^2 + 1 its just asking you to replace x with 3x - 1

OpenStudy (anonymous):

HAHA thaks so much. (: